shift summary for Eve shift 06.02.99

Jim Graham (jagraham@hep.uchicago.edu)
Thu, 03 Jun 1999 00:11:11 -0500

Shift Summary for Eve shift, 06.02.99

On Shift: JAG, AT, plus a cast of tens

Hours of beam delivered at all intensities: ~6
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 1

Accounting of lost time:

========================================================================
Runs ended this shift:
13302 - beam_832_ecal (no useful data)
13305 - long laser scan

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift:

15:30 -- Ed, Elizabeth, Hogan on access to pull regenerator out of beam
and to look at DCs.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 - Rick realized that the shadow absorber was out
while the reg was in --> previous rates and accidental
profiles way out of nominal
17:25 - SET tweaks the Reg HV
17:25 - Rick remakes the L3_driver_832 to fix L2 eff. monitoring
13302 17:41 - Run initialization hangs on time out from the
kumquats. We aren't sure if this came up as a red
button in Run Control. We found the problem in
~da_ktev/dest/target_unix_atev1_06_02.log with the
cryptic message telling us to reset the fscc controller
on fastbus crate 2. This seemed to work.
18:15 - ddd veto 1 & 2 hangs the end of the run. We page in
Margaret Votava, who increases the number of buffers
allocated in the dis database. This seems to work. But
she also points out some possible problems with l3driver
which experts are currently looking at.
18:46 - Rick Ford begins studies to test our SEM and seeds.
We put in the beam stop in anticipation of high rate.
In doing these tests, MCR finds a problem with beam
aborts, and can't raise the intensity to do the test.
13305 22:45 - while waiting for MCR, we do a long laser scan.
23:50 - Rick Ford has finished his tests up to 1E12 and things
look good. He plans to fine tune things, and send beam
to ME dump before opening the beam stop